OpenStudy (anonymous):

Okay, so I have to see if I did this correctly(: Use the given equations to find the missing domain and range elements: y=-3x+5 Domain Range x y --------------- -3 0 7 1 4 So I got: -3 10 0 -5.5 ? 7 1 4 4 2.5

OpenStudy (anonymous):

if x = 1, y = -3 + 5 = 2. The line 1 4 should be 1 2

OpenStudy (anonymous):

Oh my gosh! I put the wrong equation up! SORRY! Its \[x^2-2y=11\]

OpenStudy (anonymous):

if x = 4, y = -3*4 + 5 = -12 + 5 = -7 line 4 2.5 should be 4 -7 Try to find the value of x that gives y = 7 :]

OpenStudy (anonymous):

wow, it changes everything...

OpenStudy (anonymous):

x = -3, you will have: (-3)² - 2y = 11 9 - 2y = 11 9 - 11 = 2y y = -1

OpenStudy (anonymous):

x = 0 it's correct. for x = 1 1² - 2y = 11 1 - 2y = 11 1 - 11 = 2y -10 = 2y y = - 5 x = 4 its correct.

OpenStudy (anonymous):

With y = 7, x² - 14 = 11 x² = 25 x = 5

OpenStudy (anonymous):

So: -3 -1 0 -5.5 5 7 1 -5 4 2.5

OpenStudy (anonymous):

Yea!

OpenStudy (anonymous):

Thanx so much!!! Much Apreciated!! GBU!! :)

OpenStudy (anonymous):

Oh, a problem!! In the line 5 7 it's +- 5.

OpenStudy (anonymous):

x² = 25 gives the solution x = 5 and x = -5

OpenStudy (anonymous):

haha(: thanx!

OpenStudy (anonymous):

so instead of 4 and 2.5 its 5 and -5

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

In the row: 5 7 (x = 5, y = 7) Its correct, however, -5 7 is correct too.

OpenStudy (anonymous):

Because you have to find the value of x that gives you y = 7, and x = 5 and x=-5 gives this. Just try: x² - 2y = 11 (5)² - 2y = 11 25 - 2y = 11 14 = 2y y= 7 (-5)² - 2y = 11 25 - 2y = 11 14 = 2y y = 7

OpenStudy (anonymous):

Haha, thank!